The claim is that for 12 AM body​ temperatures, the mean is mugreater than98.6degreesF. The sample size is n=7 and the test statistic is t=2.377.
Answer #1

Solution :

=

=

=

n =

This is the right tailed test .

The null and alternative hypothesis is ,

H0 :    = 98.6

Ha : > 98.6

n =7

df = n- 1 = 7-1 = 6

t = 2.377

p value = p(t(6) > 2.377) =0.0275

P-value =0.0275